OpenStudy (anonymous):

log base 37 6 rounded to three decimal places?

OpenStudy (anonymous):

\[\left( \ln 6 \right) \div \left( \ln 37 \right) = 0.496206092\]

OpenStudy (anonymous):

thank you soo much ! im horrible at logs =/